# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

### 10 Jan 2024 v1

@@ 4.1. Dataset Creation and Curation
A validation dataset was created to assess the quality of the analyses performed by INACIA. This dataset consists of a fully annotated table with the expected results for each pipeline step. In this study, we only worked with closed cases. This means the document of interest for evaluation (the decision document called instruction) was already available for all cases. This document is a searchable PDF written by TCU staff with expertise on the case subject. It is written in natural language and divided into five main sections: basic information, claims, and requests, admissibility examination, precautionary measure analysis (periculum in mora and fumus boni iuris, and recommendations. We used instructions from closed cases to build the validation dataset.

### 19 Jan 2024 v2

@@ 4.1. Dataset Creation and Curation
A validation dataset was created to assess the quality of the analyses performed by INACIA. This dataset consists of a fully annotated table with the expected results for each pipeline step. In this study, we only worked with closed cases. This means the document of interest for evaluation (the initial instruction document) was already available for all cases. This document is a searchable PDF written by TCU staff with expertise on the case subject. It is written in natural language and divided into five main sections: basic information, claims, and requests, admissibility examination, precautionary measure analysis (periculum in mora and fumus boni iuris, and recommendations. We used instructions from closed cases to build the validation dataset.
@@ 5. Results
Although the INACIA pipeline does not always produce the expected results, it is possible to follow the reasoning of the INACIA pipeline to understand the decisions made by the agent. This helps the auditor to find the evidence. For example, the INACIA pipeline may not classify the case as admissible regarding some criterion in the admissibility examination. However, it is possible to follow the reasoning of the INACIA pipeline to understand why the case was not classified as admissible. This helps the auditor to find the evidence that supports the decision made by INACIA.